On 1 January 2017, Abracadabra Bhd entered into a 5-year franchise agreement with Yoohuu Bhd to operate a health club and spa in Pulau Lang Tengah. At that time, the company has to pay RM500,000 for rights to use the franchisor’s logo and technique. Besides, Abracadabra Bhd paid Yoohuu Bhd RM1,500,000 for special equipment to be used in the health club and spa. The equipment can be purchased separately by the holders of the franchise only. The equipment has a useful life of 5 years.

Required:

Explain briefly the accounting treatment for the year ended 31 December 2017 in each of the above situations. Justify your answer with appropriate journal entries.
Answer #1

Franchise has special accounting treatment and has different accounting concepts.

The Initial Franchise Fees paid by the Franchisee is treated as an intangible asset and it is amortized over the years for the period of franchise. The following are the Journal entries for it.

Date Particulars Debit (RM) Credit (RM)
1.1.2017 Franchise a/c Dr 500,000
To Cash a/c 500,000
31.12.2017 Franchise Fees Amortization a/c Dr 100,000
To Franchise a/c 100,000

The First Entry is made on 1st jan 2017 to record the Franchise Fees paid by Abracadabra Bhd to Yoohuu Bhd.

The Second entry is made on 31st dec 2017 to Amortize the Franchise Fees for the first year. As the Franchise is for 5 years we have divided the Franchise Fees by 5 to find out the amount to be amortized per year.

Franchise Fees to be Amortized = 500,000 / 5 = 100,000

Next is the special equipment purchased for the Franchise. The equipment is an asset and the following entries are made in the Journal.

Date Particulars Debit (RM) Credit (RM)
1.1.2017 Special Equipment Dr 1,500,000
To Cash a/c 1,500,000
31.12.2017 Depreciation a/c Dr 300,000
To Special Equipment a/c 300,000

The First Entry is made on 1st jan 2017 to record the purchase of special equipment by Abracadabra Bhd.

The Second entry is made on 31st dec 2017 to depreciate the equipment. As the useful life of the equipment is for 5 years we have divided the cost of equipment by 5 to find out the amount of depreciation.

Depreciation = 1,500,000 / 5 = 300,000